# Viscoplasticity can stabilise liquid collar motion on vertical cylinders

### v2

@@ III.1 Linear stability analysis
In the initial state, the fluid is yielded in $0\leq y\leq Y=1-\mathcal{J}$, and unyielded in $Y=1-\mathcal{J}<y\leq 1$ where the axial velocity equals the (pseudo-)plug velocity, $w=w_{\mathrm{p}}$, as defined in (21). The shear stress (17) is initially zero at the free surface, so we expect that a finite-amplitude perturbation to the stress may be required to trigger unstable evolution of the free surface. However, when $\mathcal{J}<1$, the thin-film theory predicts that there can be a linear instability given an infinitesimal perturbation to the free-surface height. To show this, when $\mathcal{J}<1$, we let
$h=1+\tilde{h}(z,t)\quad\mbox{where}\quad|\tilde{h}|\ll 1,$ (25)
for some function $\tilde{h}$. Then, from (21), the pseudo-plug velocity is
$w_{p}\sim-\frac{1}{2}B(1-\mathcal{J})^{2}-\frac{1}{2}(1-\mathcal{J})\left[2B\tilde{h}-(1+\mathcal{J})(\tilde{h}_{z}+\tilde{h}_{zzz})\right]+O(\tilde{h}^{2})\quad\mbox{when}\quad\mathcal{J}<1.$ (26)
From (26), it can be seen that, in general, the axial velocity is a function of $z$ when $|\tilde{h}|\ll 1$ and $\tilde{h}$ is not a constant function, and so there must be a genuine pseudo-plug, where the fluid is weakly yielded, in $Y<y\leq h$ [36]. In short, when $\mathcal{J}<1$, an infinitesimal perturbation to the free surface height can induce a sufficiently large increase in the free-surface stress to yield the layer and potentially trigger instability. When there is no gravity [15, 16], or when $\mathcal{J}>1$, a finite-amplitude perturbation to the free surface height is required to yield the layer and trigger instability. Balmforth & Liu [34] have previously identified and discussed how infinitesimal perturbations to the height of a thin viscoplastic film can trigger instability, even when the initial state has a plug or pseudo-plug at the free surface, in the context of inclined plane flow.
We now suppose that $\mathcal{J}<1$ and
$\tilde{h}(z,t)=Ae^{ikz+st},$ (27)
where $|A|\ll 1$. Then (22) gives the dispersion relation,
@@
The Newtonian growth rate [22] is recovered by setting $\mathcal{J}=0$ in (28). The growth rate, $\mathrm{Re}(s)$ in (28), decreases monotonically as $\mathcal{J}$ is increased, and approaches zero as $\mathcal{J}\rightarrow 1$. The deviation from the Newtonian growth rate in (28) is proportional to $\mathcal{J}^{3}$, so is small unless $\mathcal{J}\approx 1$; viscoplasticity has a minor impact on the growth rate unless the yield stress is large enough that the fluid is almost entirely unyielded in the initial state. As in the Newtonian problem [22], the system is unstable to disturbances with wavenumber $10<k<$ 1, and the most unstable wavelength corresponds to $k=k_{\mathrm{m}}\equiv\sqrt{2}/2$.
@@ III.2 Nonlinear evolution
Figure 2: Numerical solution of (22) with $B=0.1$, $J=0.05$, in a periodic domain of length $\mathcal{L}=2\pi/k_{\mathrm{m}}=2\sqrt{2}\pi$. (a) Snapshots of the evolving layer. Lines show $h(z,t)$ (blue) and $Y(z,t)$ (red); colour map illustrates the axial velocity. Arrow in upper-left panel indicates the direction of gravity. (b) Time evolution of the maximum layer height minus the average layer height (black), with the predicted exponential growth rate from linear theory (28) (dashed magenta). (c) Location, $z_{\mathrm{max}}$, of the maximum of $h(z,t)$. The leading-order asymptotic prediction (78) for the steady collar speed, $B^{3/2}|\mathcal{U}|$, is also illustrated in (c), where we have used the approximation $V=\mathcal{L}$ to calculate $\mathcal{U}$ in (78).
To analyse the nonlinear evolution, we solve (22) numerically, in a periodic domain, $0\leq z\leq\mathcal{L}$, with initial condition (23) replaced by
@@
These numerical simulations in fixed periodic domains can capture both the initial growth of the instability from a near-flat layer, leading to collar formation, and the subsequent translation of collars through the domain. In all numerical simulations presented, we set $\mathcal{L}=2\pi/k_{\mathrm{m}}=2\pi\sqrt{2}$, which means that the domain can accommodate one unstable mode, the most unstable mode from the linear stability analysis. We have also run simulations in domains of different lengths, finding qualitatively similar results in general, except that when the domain is made significantly long more than one collar may develop. When $\mathcal{L}>2n\pi$, the domain can accommodate $n$ linearly unstable modes, and we will see below that collars have length of approximately $2\pi$ when $B$ is small, so we expect that up to $n$ collars may develop in a domain of length $2n\pi<\mathcal{L}<2(n+1)\pi$. We focus on the emergence and evolution of a single collar, and so choose $\mathcal{L}=2\pi/k_{\mathrm{m}}$, which satisfies $2\pi<\mathcal{L}<4\pi$.
@@
Figure 2 shows output from a typical numerical simulation where instability and collar formation occur. In the initial state (figure 2a, $t=0$), $Y\approx 1-\mathcal{J}$ so there is flow in the layer but the velocity is relatively low throughout. After a short time ($t=40$), the free surface has become non-uniform and then ($t=80$) an emerging fluid collar is evident, which is translating in the negative $z$-direction. The growth rate of the peak height of the layer is well-approximated by the exponential growth rate predicted by linear stability analysis (figure 2b). As the collar is growing, the axial velocity, $w$, is largest in the rear half of the collar (figure 2a, $t=80$), where capillary forces and gravity are both driving fluid in the negative $z$-direction, whilst at the front of the collar, the velocity is weaker since capillary forces are acting in the opposite direction to gravity. At $t=80$, there are two points in the domain at which $Y\approx 0$; these are points at which the direction of flow changes, indicating that at the front of the emerging collar there is a weak flow in the positive $z$-direction. Subsequently, the collar gradually grows, and a thin film develops in the rest of the domain. By $t=600$, the collar has essentially stopped growing (figure 2b), and is translating at an almost constant speed (figure 2c). From figure 2(a), at $t=600$, there is a short region within the thin film, at around $z\approx 6$, where $Y=0$ and $w=0$. For larger values of $B$, the film between the front and back of the collar would be thicker and more likely to be yielded throughout, but for the relatively small values of $B$ that we focus on, it is typical that there is an unyielded region, as in figure 2(a) at $t=600$. At the front edge of the collar, $h$ undulates and there are several bumps in $Y$ (figure 2a, $t=200,600$), indicating that the direction of flow is changing repeatedly, although the velocity is low throughout this region. Predicting the late-time shape of $h$ and $Y$ is the focus of the modelling in §IV, and in §V we compare predictions from the asymptotic model to late-time output from numerical simulations.
@@ IV Small-$B$ model for quasi-steadily translating collars
In this section, we use matched asymptotics to derive a model for the translation of a slender collar of liquid down the tube wall, considering (22) in the limit $B\ll 1$. This analysis has some key differences from that of the Newtonian analogue [4]. We consider an isolated collar of liquid translating with dimensionless speed $U(t)$, with a film of uniform thickness, $h_{\infty}^{-}$, ahead of the collar, and a film of uniform thickness, $h_{\infty}^{+}$, deposited behind the collar. We no longer assume that the domain is periodic, as we did in §III, in order to describe general quasi-steady motion. We assume that $h=O(1)$ in the main body of the collar, but that the films ahead and behind are much thinner, so that $h_{\infty}^{\pm}\ll 1$. We will consider the regime in which $B$, as defined in (14), is of the same order of magnitude as the film thickness ahead of the collar, so $B=O(h_{\infty}^{-})$, and in which the deposited film behind the collar is also of the same order of magnitude, $h_{\infty}^{+}=O(B)$, although we do not, in general, enforce that $h_{\infty}^{+}=h_{\infty}^{-}$. We write $h_{\infty}^{-}=BH_{\infty}^{-}$ and $h_{\infty}^{+}=BH_{\infty}^{+}$, where $H_{\infty}^{\pm}=O(1)$, and treat $B$ as the small parameter in the problem. We could equally treat $h_{\infty}^{-}$ as the small parameter and treat $B$ as a function of $h_{\infty}^{-}$ (e.g., [4]).
@@
We consider the system in a frame of reference moving with the velocity of the collar, so that the tube is translating with speed $-U\bm{\hat{z}}$ in the moving frame. Defining $\zeta$ to be the axial coordinate in the moving frame, (22) is modified to give
@@
$h_{\zeta}+h_{\zeta\zeta\zeta}=0.$ (34)
We expect this capillary-static solution to be perfectly wetting on the tube wall, so we enforce boundary conditions, $h(0)=h_{\zeta}(0)=0$, where we have chosen the front edge of the collar, which can be placed at any arbitrary point in $\zeta$, to lie at $\zeta=0$. Then, the capillary-static solution is a small-amplitude version of an unduloid,
@@
In the Newtonian problem [4], the solution for $h$ has a three-region structure: the main body of the collar is connected to two short ‘Bretherton’ regions of length ${O}(B^{1/2})$, one at the front and one at the rear of the collar, which then match onto the films of uniform thickness ahead of, and behind, the translating body. We name these ‘Bretherton’ regions given their similarity to analogous regions in the solution for the motion of long bubbles in liquid-filled tubes [43, 44]. We will show that, in the limit $B\ll 1$, similar asymptotic regions exist in the viscoplastic problem as in the Newtonian problem. However, multiple additional regions emerge, which must be considered in order to construct complete solutions for $h$ and $Y$. Figure 3 illustrates the locations and sizes of the different asymptotic regions. Beyond the Bretherton regions, in which the fluid turns out to be almost entirely yielded with $Y\approx h$ almost everywhere, there are ‘intermediate regions’, with characteristic axial length scale of size $O(B^{1/4})$, between the main collar body and the Bretherton regions, in which $Y$ transitions from being close to $h$ to being small compared to $h$. There are also very short inner regions, of length $O(B)$, within the Bretherton region at the front of the collar (figure 3b), and there are ‘rigidification regions’, which have characteristic length scale of size $O(B^{1/2})$, between the Bretherton region and the uniform film, in which the fluid transitions from being fully yielded to being unyielded; we leave detailed discussion of the inner regions within the front Bretherton region, and of the rigidification regions, to appendices B and C, respectively. Since the fluid in the uniform films ahead of and behind the collar must be unyielded, there must be two points, call them $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$, such that $Y\rightarrow 0^{+}$ as $\zeta\rightarrow\zeta_{0}^{-}$, $\zeta\rightarrow\zeta_{1}^{+}$, and $Y=0$ for $\zeta<\zeta_{1}$ or $\zeta>\zeta_{0}$. Since each of the regions at the edges of the main collar body is asymptotically short, we expect that $|\zeta_{1}|\ll 1$ and $|\zeta_{0}-2\pi|\ll 1$. The two rigidification regions are in neighbourhoods of the points $\zeta=\zeta_{1}$ and $\zeta=\zeta_{0}$, and we show in appendix C below that the Bretherton regions sit at a distance of approximately $O(B^{1/2}\log(1/B))$ from $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$, respectively (figure 3b,c).
@@ IV.2 Main collar body
$h=h_{0}(\zeta)+B\log Bh_{1}(\zeta)+Bh_{2}(\zeta)+\dots,\quad Y=B^{1/4}Y_{0}(\zeta)+B^{1/2}Y_{1}(\zeta)+\dots.$ ($)39𝑎𝑏$
Given the expansions (($)39𝑎𝑏$), (30) and (31) imply
[…]
@@
[…]
@@
[…]
@@
$h\sim\frac{V}{4\pi}\zeta^{2}+B\log{B}\,a_{1}\zeta+B\left(a_{2}+\frac{4\pi\mathcal{J}}{V}\zeta\log{\zeta}+\zeta\left[b_{2}+1-\frac{2\pi\mathcal{J}}{V}(1+2\log{2})\right]\right)\quad\mbox{as}\quad\zeta\rightarrow 0^{+},$ (49)

### v3

@@ III.1 Instability of an almost flat layer
For the uniform film, the fluid is fully yielded in $0\leq y\leq Y=1-\mathcal{J}$.
When $\mathcal{J}=0$, a uniform film is known to be linearly unstable to long-wave perturbations [22]. When $0<\mathcal{J}<1$, if the layer height is exactly uniform, we expect there to be an unyielded plug adjacent to the free surface. For unstable growth to occur, the perturbation applied to the layer must be strong enough to yield the fluid at interface. A small perturbation to the stress would not, in general, be sufficient to break the plug and induce instability, but a perturbation to the velocity field could raise the stress above the yield stress and yield the layer, as discussed by Balmforth & Liu [44] for the related problem of inclined plane flow. Whether there is a plug at the interface after a perturbation is applied to $h$ may depend on the amplitude of that perturbation.
The thin-film theory employed in §II assumes that pseudo-plugs exist wherever the magnitude of the shear stress drops below the yield stress, and that these pseudo-plugs are able to deform. Liu et al. [45] demonstrated that, for gravity currents flowing down inclines, where the free-surface height deviates from uniform by an amount of the order of $\epsilon$, the thin-film parameter, then a solution can exist with a genuinely rigid plug at the free surface. An alternative thin-film theory with a rigid plug at the interface is then required, which is combined with standard thin-film theory in regions where the free-surface height deviates from uniform by more than $O(\epsilon)$ [45]. Analogously, in the problem we consider here, we expect that the perturbation applied to the free-surface height initially must be large enough that the fluid adjacent to the free surface is a marginally-yielded pseudo-plug, for the thin-film evolution equation (22) to hold. In the following, we assume that the layer is perturbed in such a way that the fluid at the interface is yielded and (22) holds. As in [45], an alternative theory may be required to investigate the behaviour of a flat layer subjected to perturbations with amplitude of $O(\epsilon)$ or smaller, taking into account the possibility of a rigid plug at the interface; we do not pursue this here, instead focusing on the finite-amplitude instability and dynamics described by (22). Viscoplastic thin-film theory is also known to break down in asymptotically small regions around points where the direction of flow changes [36], so it is possible that small rigid plugs may also exist in these regions that are not predicted by the theory; we do not examine these asymptotically small regions in detail, but assume that the thin-film theory still provides a good approximation to the global dyanmics.
To examine the instability of an almost flat layer, we assume that $\mathcal{J}<1$ and consider a sinusoidally perturbed free surface,
$h(z,t)=1+Ae^{ikz+st},$ (25)
where $\epsilon\ll|A|\ll 1$. Substituting (25) into (22) gives the perturbed surface separating the fully-yielded region and pseudo-plug,
$Y\sim 1-\mathcal{J}+\left[1-\frac{\mathcal{J}ik(1-k^{2})}{B}\right]Ae^{ikz+st}+O(|A|^{2}),$ (26)
and the pseudo-plug velocity,
$w_{\mathrm{p}}\sim-\frac{B}{2}(1-\mathcal{J})^{2}-\frac{1}{2}(1-\mathcal{J})\left[2B-{i}k(1+\mathcal{J})(1-k^{2})\right]Ae^{ikz+st}+O(|A|^{2}).$ (27)
After linearising, (22) gives
@@
The Newtonian growth rate [22] is recovered by setting $\mathcal{J}=0$ in (28). The growth rate, $\mathrm{Re}(s)$ in (28), decreases monotonically as $\mathcal{J}$ is increased, and approaches zero as $\mathcal{J}\rightarrow 1$. The deviation from the Newtonian growth rate in (28) is proportional to $\mathcal{J}^{3}$, so is small unless $\mathcal{J}\approx 1$; viscoplasticity has a minor impact on the growth rate unless the yield stress is large enough that the fluid is almost entirely unyielded in the initial state. As in the Newtonian problem [22], the system is unstable to disturbances with wavenumber $0<k<1$, and the most unstable wavelength corresponds to $k=k_{\mathrm{m}}\equiv\sqrt{2}/2$, although, as discussed above, this argument holds in the viscoplastic case only when the free-surface perturbations are large enough to yield the fluid at the interface.
@@ III.2 Nonlinear evolution
Figure 2: Numerical solution of (22) with $B=0.1$, $J=0.05$, in a periodic domain of length $\mathcal{L}=2\pi/k_{\mathrm{m}}=2\sqrt{2}\pi$. (a) Snapshots of the evolving layer. Lines show $h(z,t)$ (blue) and $Y(z,t)$ (red); colour map illustrates the axial velocity. Arrow in upper-left panel indicates the direction of gravity. (b) Time evolution of $\max_{z}(h)-1$ (black) with the predicted early-time exponential growth rate from (28) (dashed magenta), and $\min_{z}(h)$ (blue). (c) Location, $z_{\mathrm{max}}$, of the maximum of $h(z,t)$. The leading-order asymptotic prediction (78) for the steady collar speed, $B^{3/2}|\mathcal{U}|$, is also illustrated in (c), where we have used the approximation $V=\mathcal{L}$ to calculate $\mathcal{U}$ in (78).
To analyse the nonlinear evolution of an initially almost flat layer, we solve (22) numerically in a periodic domain, $0\leq z\leq\mathcal{L}$, with initial condition (23) replaced by
@@
These numerical simulations in fixed periodic domains can capture both the initial growth of the instability from a near-flat layer, leading to collar formation, and the subsequent translation of collars through the domain. In all numerical simulations presented, we set $\mathcal{L}=2\pi/k_{\mathrm{m}}=2\pi\sqrt{2}$, which means that the domain can accommodate one unstable mode, the most unstable mode from (28). We have also run simulations in domains of different lengths, finding qualitatively similar results in general, except that when the domain is made significantly long more than one collar may develop. We will see below that collars have length of approximately $2\pi$ when $B$ is relatively small, so we expect that up to $n$ collars may develop in a domain of length $2n\pi<\mathcal{L}<2(n+1)\pi$. We focus on the emergence and evolution of a single collar, and so choose $\mathcal{L}=2\pi/k_{\mathrm{m}}$, which satisfies $2\pi<\mathcal{L}<4\pi$. It may be possible that in longer domains, multiple collars may interact and there may be secondary instabilities; we do not pursue this here, instead focusing on the dynamics of individual collars.
@@
Figure 2 shows output from a typical numerical simulation where instability and collar formation occur. In the initial state (figure 2a, $t=0$), $Y\approx 1-\mathcal{J}$ so there is flow in the layer but the velocity is relatively low throughout. After a short time ($t=40$), the free surface has become non-uniform and then ($t=80$) an emerging fluid collar is evident, which is translating in the negative $z$-direction. The growth rate of the peak height of the layer is well-approximated by the exponential growth rate predicted in (28) (figure 2b). As the collar is growing, the axial velocity, $w$, is largest in the rear half of the collar (figure 2a, $t=80$), where capillary forces and gravity are both driving fluid in the negative $z$-direction, whilst at the front of the collar, the velocity is weaker since capillary forces are acting in the opposite direction to gravity. At $t=80$, there are two points in the domain at which $Y\approx 0$; these are points at which the direction of flow changes, indicating that at the front of the emerging collar there is a weak flow in the positive $z$-direction. Subsequently, the collar gradually grows, and a thin film develops in the rest of the domain.
By $t=600$, the collar has essentially stopped growing (figure 2b), and is translating at an almost constant speed (figure 2c). From figure 2(a), at $t=600$, there is a short region within the thin film, at around $z\approx 6$, where $Y=0$ and $w=0$. For larger values of $B$, the film between the front and back of the collar would be thicker and more likely to be yielded throughout, but for the relatively small values of $B$ that we focus on, it is typical that there is an unyielded region, as in figure 2(a) at $t=600$. The minimum layer thickness approaches a non-zero value at late times (figure 2b), confirming that a liquid film is continually deposited behind the collar. At the front edge of the collar, $h$ undulates and there are several bumps in $Y$ (figure 2a, $t=200,600$), indicating that the direction of flow is changing repeatedly, although the velocity is low throughout this region. Predicting the late-time shape of $h$ and $Y$ is the focus of the modelling in §IV, and in §V we compare predictions from the asymptotic model to late-time output from numerical simulations.
@@ IV Small-$B$ model for quasi-steadily translating collars
In this section, we use matched asymptotics to derive a model for the translation of a slender collar of liquid down the tube wall, considering (22) in the limit $B\ll 1$. Focusing on the limit $B\ll 1$ enables us to develop an asymptotic model that will provide a detailed picture of the structure of a translating collar. We will subsequently compare results to numerical solutions of the evolution equation (22), which is also valid when $B$ is not small. The analysis in this section has some key differences from that of the Newtonian analogue [5]. We consider an isolated collar of liquid translating with dimensionless speed $U(t)$, with a film of uniform thickness, $h_{\infty}^{-}$, ahead of the collar, and a film of uniform thickness, $h_{\infty}^{+}$, deposited behind the collar. We no longer assume that the domain is periodic, as we did in §III, in order to describe general quasi-steady motion. We assume that $h=O(1)$ in the main body of the collar, but that the films ahead and behind are much thinner, so that $h_{\infty}^{\pm}\ll 1$. We will consider the regime in which $B$, as defined in (14), is of the same order of magnitude as the film thickness ahead of the collar, so $B=O(h_{\infty}^{-})$, and in which the deposited film behind the collar is also of the same order of magnitude, $h_{\infty}^{+}=O(B)$, although we do not, in general, enforce that $h_{\infty}^{+}=h_{\infty}^{-}$. We write $h_{\infty}^{-}=BH_{\infty}^{-}$ and $h_{\infty}^{+}=BH_{\infty}^{+}$, where $H_{\infty}^{\pm}=O(1)$, and treat $B$ as the small parameter in the problem. We could equally treat $h_{\infty}^{-}$ as the small parameter and treat $B$ as a function of $h_{\infty}^{-}$ (e.g., [5]).
@@
We consider the system in a frame of reference moving with the velocity of the collar, so that the tube is translating with speed $U\bm{\hat{z}}$ in the moving frame. Defining $\zeta$ to be the axial coordinate in the moving frame, (22) is modified to give
@@
$h_{\zeta}+h_{\zeta\zeta\zeta}=0,$ (34)
which is the same whether the fluid is Newtonian or viscoplastic. We expect this capillary-static solution to be perfectly wetting on the tube wall, so we enforce boundary conditions, $h(0)=h_{\zeta}(0)=0$, where we have chosen the front edge of the collar, which can be placed at any arbitrary point in $\zeta$, to lie at $\zeta=0$. Then, the capillary-static solution is a small-amplitude version of an unduloid,
@@
In the Newtonian problem [5], the solution for $h$ has a three-region structure: the main body of the collar is connected to two short ‘Bretherton’ regions of length ${O}(B^{1/2})$, one at the front and one at the rear of the collar, which then match onto the films of uniform thickness ahead of, and behind, the translating body. We name these ‘Bretherton’ regions given their similarity to analogous regions in the solution for the motion of long bubbles in liquid-filled tubes [46, 47]. We will show that, in the limit $B\ll 1$, similar asymptotic regions exist in the viscoplastic problem as in the Newtonian problem. However, multiple additional regions emerge, which must be considered in order to construct complete solutions for $h$ and $Y$. Figure 3 illustrates the locations and sizes of the different asymptotic regions. Beside the Bretherton regions, in which the fluid turns out to be almost entirely yielded with $Y\approx h$ almost everywhere, there are ‘intermediate regions’, with characteristic axial length scale of size $O(B^{1/4})$, between the main collar body and the Bretherton regions, in which $Y$ transitions from being close to $h$ to being small compared to $h$. There are also very short inner regions, of length $O(B)$, within the Bretherton region at the front of the collar (figure 3b), and there are ‘rigidification regions’, which have characteristic length scale of size $O(B^{1/2})$, between the Bretherton region and the uniform film, in which the fluid transitions from being fully yielded to being unyielded; we leave detailed discussion of the inner regions within the front Bretherton region, and of the rigidification regions, to appendices B and C, respectively. Since the fluid in the uniform films ahead of and behind the collar must be unyielded, there must be two points, call them $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$, such that $Y\rightarrow 0^{+}$ as $\zeta\rightarrow\zeta_{0}^{-}$, $\zeta\rightarrow\zeta_{1}^{+}$, and $Y=0$ for $\zeta<\zeta_{1}$ or $\zeta>\zeta_{0}$. Since each of the regions at the edges of the main collar body is asymptotically short, we expect that $|\zeta_{1}|\ll 1$ and $|\zeta_{0}-2\pi|\ll 1$. The two rigidification regions are in neighbourhoods of the points $\zeta=\zeta_{1}$ and $\zeta=\zeta_{0}$, and we show in appendix C below that the Bretherton regions sit at a distance of approximately $O(B^{1/2}\log(1/B))$ from $\zeta=\zeta_{0}$ and $\zeta=\zeta_{1}$, respectively (figure 3b,c).
@@ IV.2 Main collar body
$h=h_{0}(\zeta)+B\log Bh_{1}(\zeta)+Bh_{2}(\zeta)+\dots,\quad Y=B^{1/4}Y_{0}(\zeta)+B^{1/2}Y_{1}(\zeta)+\dots.$ (39a,b)
Given the expansions ((39a,b)), (30) and (31) imply
$h_{0}^{\prime}+h_{0}^{\prime\prime\prime}=0,\quad h_{1}^{\prime}+h_{1}^{\prime\prime\prime}=0,\quad\frac{1}{2}(1-h_{2}^{\prime}-h_{2}^{\prime\prime\prime})Y_{0}^{2}+\mathcal{U}=0,$ (40a,b,c)
@@
In (41), we have assumed that $1-h_{2}^{\prime}-h_{2}^{\prime\prime\prime}\geq 0$, which follows from ((40a,b,$c)$ since we expect $\mathcal{U}\leq 0$, and we have assumed that $Y>0$.
The solution to ((40a,b,c)a) is the capillary-static solution (35), namely,
@@
for some constants $a_{1}$, $a_{2}$ and $b_{2}$ that will be determined below. The final three terms in (46) capture a weak gravity-induced tilting of the collar. Finally, ((40a,b,$c)$ and (41) give
@@
$h\sim\frac{V}{4\pi}\zeta^{2}+B\log{B}\,a_{1}\zeta+B\left(a_{2}+\frac{4\pi\mathcal{J}}{V}\zeta\log{\zeta}+\zeta\left[b_{2}+1-\frac{2\pi\mathcal{J}}{V}(1+2\log{2})\right]\right)~\mbox{as}~\zeta\rightarrow 0^{+},$ (49)